# On the first test of the Weak Equivalence Principle in low Earth orbit

### v4

@@ III Forces acting on Microscope accelerometers and their control
For such biases orders of magnitude above the target not to impair the measurement, it is mandatory that they remain constant. This will not be the case in the presence of patch charges on the surface of the TMs and their electrodes. Differences in the microscopic crystal structures, and impurities, over the surface of the test masses and their surrounding electrodes give rise to local potential variations and hence to spurious forces and torques (patch effects). The smaller the gap between the surfaces, the larger the effect. Random fluctuations of these potentials translate into a $1/\sqrt{\nu}$ spectral density. Moreover, occasional dislocations of these microstructures, whatever the cause, may produce sudden variations of the mean value of the fluctuations (Sec. IV).
@@
Small force experiments require to take care of thermal effects. A well known “killer” in testing the WEP is the radiometer effect. Residual air pressure in the cage enclosing the TM, combined with a non zero temperature gradient along the symmetry axis, gives rise to a spurious acceleration along this axis which varies with the synodic frequency of the spacecraft relative to the Earth, it is inevitably different for the two masses, and therefore competes directly with a violation signal should the symmetry axis be –like in Microscope– the sensitive axis of the WEP test. The radiometer effect has been a matter of concern20 but direct in orbit measurements have shown that a very good thermal stability has been achieved (mostly by thermal insulation of the spacecraft), such that, at the $10^{-15}$ target level of Microscope, systematic thermal effects are smaller than random errors (SUEP) or close to them (SUREF).9; 8
IV Limiting noise, patch charges and unstable zero
@@
Ever since the preliminary Microscope results,5 the residual random acceleration differences in the frequency region of the signal were attributed to thermal noise from internal damping in the gold wires.
@@
In the case of internal damping the line which best fits the spectral density of the limiting noise decreases with the frequency as $1/\sqrt{\nu}$ and reads:21
$S_{a}^{{}^{1/2}}(\nu)=f_{{}_{TkQ}}\mdot\frac{1}{\sqrt{2\pi\nu}}$ (4)
@@
$f_{{}_{TkQ}}=\sqrt{4k_{{}_{B}}T\mdot\left(\frac{k_{1}}{m_{1}^{2}}+\frac{k_{2}}{m_{2}^{2}}\right)\mdot\frac{1}{Q}}\ \ \ \ \ \ .$ (5)
@@
For a measurement session of duration $t_{int}$, as long as the acceleration noise is random, and therefore is bound to decrease as the square root of the integration time, the value of Eq.(4) at $\nu_{{}_{EP}}$ sets the limit $\delta$ to the violation signal established by the measurement session, according to the relationship:
$\delta\mdot g_{{}_{drive}}=S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})/\sqrt{t_{int}}$ (6)
@@
where $g_{{}_{drive}}=7.9\,\rm ms^{-2}$ is the gravitational acceleration from Earth at the spacecraft orbital distance, $\delta\cdot g_{drive}$ is a tiny acceleration difference between the TMs of the SU at the known signal frequency $\nu_{{}_{EP}}$, and $\delta$ is a solve-for parameter resulting from the confrontation between the predicted acceleration differences and their measured values (as defined in Ref.(8), Eq.(6)). [^1] Note that $\delta$ is never zero because thermal noise goes to zero only for an infinite measurement time. Should a WEP violation signal be buried by fluctuations around the best fitting line given by Eq.(4), the smaller their size, the shorter the additional integration time that would be required for it to emerge and be detected. These fluctuations yield the value $\sigma$ of the resulting $\delta\pm\sigma$ for the measurement session.
@@
Early on, the spin rate was designed to be between $3$ and $5$ times the orbital frequency $\nu_{orb}$.18 Once in orbit it was set at $\frac{9}{2}\,\nu_{orb}$ (V2 mode). In 2017 it was reported that at this spin rate the thermal noise of SUEP at the signal frequency of $0.92\,\rm{mHz}$ was about $2\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ (Fig. 3, top right plot), while it was expected to be 10 times lower (Ref.(12), Fig. 11 top plot). In order to mitigate the problem the spin rate was increased to $\frac{35}{2}\,\nu_{orb}$ (V3 mode), moving the signal frequency from $0.92\,\rm mHz$ to $3.11\,\rm mHz$, with an expected noise reduction by a factor of $1.8$. Instead, the central plot of Fig. 3 shows that the reduction was by a factor of $4$, due to the fact that in going to spin mode V3 the best fitting line given by Eq.(4) jumped downwards (together with all fluctuations around it). At the new signal frequency of $3.11\,\rm mHz$ the limiting noise is still a factor between 2 and 3 higher than expected (Ref.(12), Fig. 11 top plot).
@@
It is apparent from Eq.(4) that the vertical position of the line (and of the plotted curve) is dictated by the acceleration $f_{TkQ}$, given by Eq.(5). The parameters that this acceleration depends upon can hardly have changed, in going from V2 to V3, by the amount required to explain the extra gain factor of about 2.2, never mentioned in later publications and unexplained to this date. Indeed, this event may not be exceptional, or due to the transition from lower to higher spin rate, since we observe very similar jumps even in sequential sessions at the same spin rate.
@@
Figure 3: SUEP session #218, Slide 24 of Ref.(22). By increasing the spin frequency from mode V2 to mode V3 the frequency of the signal $\nu_{{}_{EP}}$ ($f_{{}_{EP}}$ in the slide) is moved to a higher value where noise is lower (as $1/\sqrt{\nu}$). This will ensure a noise reduction by a factor 1.8. However, it also happens that the best fitting line expressed by Eq.(4) undergoes a downward parallel translation (central plot) to a lower level of noise at all frequencies, including the new $\nu_{{}_{EP}}=3.11\,\rm mHz$, such that the total noise reduction is by a factor of 4 (unexplained).
@@
We cannot exclude a variation of the stiffness after a calibration, because in the calibration the quadratic acceleration term must be kept sufficiently small, which is achieved by offsetting the TM to a new zero, hence affecting the stiffness, $f_{{}_{TkQ}}$ and $S^{{}^{1/2}}(\nu_{{}_{EP}})$.23 Conversely, in sequential sessions, or within the same session analyzed with equally valid methods, we expect no such variations. Should any anomalies occur, they will clearly appear in the values of $Q$ because $Q\propto 1/S(\nu_{{}_{EP}})$. Moreover, the extent and the relevance of the anomaly can be assessed by comparison with $Q$ measurements carried out in ground tests of WEP as well as in experiments for the measurement of the universal constant of gravity and the detection of gravitational waves.
@@
In order to carry out this analysis we need the values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ for all the valid sessions, but there is no table listing these values in the Microscope publications, as there is for $\delta_{i}\pm\sigma_{i}$,8 although accelerations are the measured physical quantities while $\delta$ is a derived number (see Ref.(8), Eq. (6)) and moreover $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ does not depend on the integration time –unlike $\delta$. With all $\delta$ values from Ref.(8), Tables 6, 7 and the duration of each session from Ref.(24) we use Eq.(6) to derive all the corresponding values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$. The correctness of our derived values can be checked when the measured ones are published. [^2] In Ref. (15), Fig. 11, the values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ are plotted for all sessions, including those which have not contributed to the final result of the WEP test, the purpose being to demonstrate an overall stability after the first few sessions of the mission. Although the numerical values of the plotted marks are not easy to read in the vertical scale used, for the 120-orbit SUEP session #236 (V3 mode, blue curve in Fig. 11, error bar shown on V2 mark by mistake) we can see that the plotted mark is much too high (by at least a factor of 3) with respect to the best fit value of about $2\times 10^{-11}\,\rm ms^{-2}/\sqrt{Hz}$ reported in Fig. 6, Ref.(8). The discrepancy is confirmed using Eq.(6), because the plotted mark leads to a value of $\delta_{236}$ at least 3 times higher than reported in Ref. (8), Table 7. For other sessions of 120-orbit duration like #236, such as #238, #254 and #404 we know from Eq.(6) that the ratios of their $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ values to that of session #236 must be the same as the ratios of the corresponding values of $\delta$ (available in Table 7 of Ref.(8)). While these ratios are never close to 1, their individual marks are all plotted at about the same height in Fig. 11 of Ref.(15). Thus, if this figure is correct, it contradicts Table 7 of Ref.(8) as well as Fig. 1 of Ref.(4) and the final Microscope test of WEP published therein.
@@
Then, given the equilibrium temperature,8 the masses and the measured stiffnesses, the quality factor $Q$ is inferred from Eq.(5). Since $k_{1}\,,k_{2}$ may vary because of calibrations we compute also $k/Q$ assuming the same unknown $k$ for both TM’s. These quantities are listed in Table 1 for SUEP and in Table 2 for SUREF, along with the percentage $p$ of artificial data that have been introduced in each session after the elimination of glitches.
@@
Table 1: SUEP sensor unit. Session number in V2 or V3 spin mode; session $\delta$ and percentage of glitches $p$ from Tables 6,7 and 4,5 of Ref.(8); calculated values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$, $Q$, $k/Q$. $\downarrow$ indicates a jump between sequential sessions; * an $M$ vs $A$ discrepancy (see text below).
@@
Two methods of analysis, $M$-$ECM$ ($M$) and $ADAM$ ($A$), have been employed to estimate the experiment parameters, including $\delta$ and $\sigma$. Both methods use artificially reconstructed data to account for the missing data resulting from the elimination of glitches. [^3] Method $M$ estimates in the time domain the missing data, maximizing the likelihood conditional on the observed data, and then computes the least-squares estimate of the regression parameters. An estimation of the SD is also produced in the process. Method $A$ performs the parameter estimation in the frequency domain. As such it requires an uninterrupted, regularly spaced time series, which is obtained by filling the gaps left by the removal of glitches with the artificial data estimated by $M$. The two methods are stated to be equivalent, but in a few cases the values of $\delta$ calculated by $M$ and $A$ differ considerably (by a factor $\simeq 2$ in sessions #218, #438 and by a factor 7 in #442), implying that the same data stream leads to a different SD according to one or the other method. SUEP session #442 is an instructive extreme case. As reported in Ref.(8), Table 7, the noise is essentially the same but the values of $\delta$ differ by a factor 7; $\delta$ representing a level of WEP violation compatible with the measured acceleration difference noise. That is, for the same measurement session, in the same experimental conditions, the same data with the same level of noise lead to almost one order of magnitude difference in the evaluation of $\delta$. The difference is particularly evident when the results are interpreted in terms of the physics of the dominant thermal noise from internal damping (widely different values of $Q$, up to an unrealistic large value in session #442). SUREF session #778-1 is the only one, among all 32 sessions ($19$ SUEP $+$ $13$ SUREF), not affected by glitches, hence without artificial data. In this case, the results are exactly the same for $\delta$ and only slightly different for $\sigma$, as one expects if the same data set is analyzed with two equally valid methods (Ref.(8), Table 6). Sessions #442 and #778-1 show that the $M$ and $A$ methods are in agreement when applied to a time series of real measured data, but give different results (in particular different $\delta$ values) in the presence of artificial reconstructed data. Since the artificial data in $A$ are those estimated by $M$, it is the way the two methods manipulate these data that makes the difference.
Table 2: Same as Table 1 for SUREF sensor unit. A $\bullet$ has been added to indicate sessions with $Q$ values much larger than the largest value of $118$ measured in ground tests. 17
@@
In SUEP session #218 the values of $\delta$ differ by a factor of about 2 depending on the type of analysis. This session (120 orbits) was the basis of the early results,5 and was further elaborated two years later,12 ending up with a different value of the spectral density at the signal frequency. Even discarding the earlier results, it is apparent that the latest analysis cannot be considered conclusive. $Q$ values differing by a factor 3.5, for the same oscillator in the same conditions, are inexplicable.
The reference sensor SUREF behaves quite differently from SUEP. Thermal noise is about a factor $\sqrt{2}$ smaller than for SUEP (assuming the same $Q$), due to a factor $2$ smaller $(k_{1}/m_{1}^{2}+k_{2}/m_{2}^{2})$ term, and residual systematic errors are comparable to the stochastic errors.8 Four out of nine sessions have been split because of sudden jumps in the mean value of the differential acceleration that required the two data segments to be treated as distinct experiments (in SUEP only one session, #326, out of 18 has been split). Unlike glitches, these jumps do not occur on all accelerometers simultaneously, hence they cannot be attributed to the spacecraft but are likely to originate in the accelerometers themselves, and last much longer (tens of seconds). In SUREF, in addition to the three cases of anomalous jumps mentioned above, we observe three sessions with large $M$ vs $A$ discrepancies (see Table 2). The values of $Q$ are typically larger than in SUEP and in three cases much larger than the largest value of $118$ reported in ground measurements.17
Early results 5 were based for SUREF on session #176 (62 orbits). The spectral density $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ reported in Ref.(5) was confirmed in Ref.(12) and attributed to $1/\sqrt{\nu}$ thermal noise, yielding $\delta=3.75\times 10^{-15}$ and $Q=65$, consistent with ground measurements. Instead, the final values of $\delta_{{M}}$ and $\delta_{{A}}$ are much smaller (Table 2), and the corresponding $Q$’s of 317 and 283 are far too high to be realistic on the basis of ground tests. Note that session #176 has $40\%$ of reconstructed data.
@@
As shown in Sec. III, the control of the TMs requires to control large acceleration biases. The acceleration bias $a_{bias}$ and the voltage $V_{bias}$ applied to control it are related through the scale factor $s_{{}_{f}}$:
@@
$s_{{}_{f}}=\frac{2\pi\epsilon_{\circ}(2r)}{md}(V_{p}-V_{p}^{\prime})$ (8)
@@
$\delta a{{}_{fpp}}=\frac{2\pi\epsilon_{\circ}(2r)}{md}\delta V_{fpp}V_{bias}\ \ \ \ .$ (9)
Using Eqs.(7) and (8), it leads to the relationship
@@
As recalled in Sec. III, the difference of acceleration biases between the TMs of SUEP is $\simeq 10^{-6}\,\rm ms^{-2}$. According to Eq.(10), a patch fluctuating potential $\delta V_{fpp}\simeq 0.3\,\mu V$ at the V3 signal frequency $\nu_{{}_{EP}}=3.11\,\rm mHz$ would produce a spurious acceleration of $4\times 10^{-14}\,\rm ms^{-2}$, as large as the average acceleration difference measured in all SUEP sessions (once normalized by session duration) at $\nu_{{}_{EP}}$ (Ref.(15), Fig. 10) and enough to change it by $100\%$ either way.
In GOCE a source of low frequency acceleration noise with $1/\sqrt{\nu}$ dependence below $0.01\,\rm Hz$ was found to be due to fluctuating potentials generated by charge patches on the test mass and its electrodes.14 GOCE was set up as a gravity gradiometer with $0.5\,\rm m$ separation of the TMs and a measurement bandwidth from $500\,\rm mHz$ to $0.1\,\rm Hz$. At $3\,\rm mHz$ the acceleration difference sensitivity was of $2\times 10^{-12}\,\rm ms^{-2}/\sqrt{Hz}$ along the ultra sensitive $Y,Z$ axes (more than one order of magnitude better than that of SUEP sensitive axis) and $4\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ along the least sensitive $X$ axis, with a contribution from patch fluctuating potentials of about $4\times 10^{-13}\,\rm ms^{-2}/\sqrt{Hz}$ and $2\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ respectively.25,  [^4] With no wire the zero of each TM control (dictated by geometry, construction, mounting and electronics) does not coincide with the zero of other forces acting on the test mass, such as Earth tides and electrostatic stiffness. Applying a constant bias is therefore inevitable in the Microscope design, regardless of the gold wire. Hence, any fluctuation of the bias in the frequency region of the WEP violation signal is relevant when aiming at very high precision.
@@
An unstable zero is a known issue in WEP tests, related to the problem of initial conditions (or release) errors, affecting all tests not designed as intrinsically null experiments, including those with laser tracked satellites, celestial bodies and cold atoms 26; 27; 28. In other tests, primarily those based on torsion balances, the physical observable is a true null.29
V Outliers, gaps and artificial data

### v6

@@ 3 Forces acting on Microscope accelerometers and their control
For such biases orders of magnitude above the target not to impair the measurement, it is mandatory that they remain constant. This will not be the case in the presence of patch charges on the surface of the TMs and their electrodes. Differences in the microscopic crystal structures, and impurities, over the surface of the test masses and their surrounding electrodes give rise to local potential variations and hence to spurious forces and torques (patch effects). The smaller the gap between the surfaces, the larger the effect. Random fluctuations of these potentials translate into a $1/\sqrt{\nu}$ spectral density. Moreover, occasional dislocations of these microstructures, whatever the cause, may produce sudden variations of the mean value of the fluctuations (Sec. 4).
@@
Small force experiments require to take care of thermal effects. A well known “killer” in testing the WEP is the radiometer effect. Residual air pressure in the cage enclosing the TM, combined with a non zero temperature gradient along the symmetry axis, gives rise to a spurious acceleration along this axis which varies with the synodic frequency of the spacecraft relative to the Earth, it is inevitably different for the two masses and therefore competes directly with a violation signal should the symmetry axis be –like in Microscope– the sensitive axis of the WEP test. The radiometer effect has been a matter of concern (Nobili et al (2001)) but direct in orbit measurements have shown that a very good thermal stability has been achieved (mostly by thermal insulation of the spacecraft), such that, at the $10^{-15}$ target level of Microscope, systematic thermal effects are smaller than random errors (SUEP) or close to them (SUREF) (Rodrigues et al (2022b); note that the total residual systematic errors in Tables 13, 14 of this paper are incorrect. For the correct values see Tables 10, 11 in Touboul et al (2022b).
4 Limiting noise, patch charges and unstable zero
@@
Ever since the preliminary Microscope results (Touboul et al (2017)) the residual random acceleration differences in the frequency region of the signal were attributed to thermal noise from internal damping in the gold wires.
@@
In the case of internal damping the line which best fits the spectral density of the limiting noise decreases with the frequency as $1/\sqrt{\nu}$ and reads (Saulson (1990)):
$S_{a}^{{}^{1/2}}(\nu)=f_{{}_{TkQ}}\cdot\frac{1}{\sqrt{2\pi\nu}}$ (4)
@@
$f_{{}_{TkQ}}=\sqrt{4k_{{}_{B}}T\cdot\left(\frac{k_{1}}{m_{1}^{2}}+\frac{k_{2}}{m_{2}^{2}}\right)\cdot\frac{1}{Q}}\ \ \ \ \ \ .$ (5)
@@
For a measurement session of duration $t_{int}$, as long as the acceleration noise is random, and therefore is bound to decrease as the square root of the integration time, the value of Eq.(4) at $\nu_{{}_{EP}}$ sets the limit $\delta$ to the violation signal established by the measurement session, according to the relation:
$\delta\cdot g_{{}_{drive}}=S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})/\sqrt{t_{int}}$ (6)
@@
where $g_{{}_{drive}}=7.9\,\rm ms^{-2}$ is the gravitational acceleration from Earth at the spacecraft orbital distance, $\delta\cdot g_{drive}$ is a tiny acceleration difference between the TMs of the SU at the known signal frequency $\nu_{{}_{EP}}$, and $\delta$ is a solve-for parameter resulting from the confrontation between the predicted acceleration differences and their measured values (as defined in Touboul et al (2022b), Eq.(6)). [^1] Note that $\delta$ is never zero because thermal noise goes to zero only for an infinite measurement time. Should a WEP violation signal be buried by fluctuations around the best fitting line given by Eq.(4), the smaller their size, the shorter the additional integration time that would be required for it to emerge and be detected. These fluctuations yield the value $\sigma$ of the resulting $\delta\pm\sigma$ for the measurement session.
@@
Early on, the spin rate was designed to be between $3$ and $5$ times the orbital frequency $\nu_{orb}$ (Touboul et al (2012)). Once in orbit it was set at $\frac{9}{2}\,\nu_{orb}$ (V2 mode). In 2017 it was reported that at this spin rate the thermal noise of SUEP at the signal frequency of $0.92\,\rm{mHz}$ was about $2\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ (Fig. 3, top right plot), while it was expected to be 10 times lower (Touboul et al (2019), Fig. 11 top plot). In order to mitigate the problem the spin rate was increased to $\frac{35}{2}\,\nu_{orb}$ (V3 mode), moving the signal frequency from $0.92\,\rm mHz$ to $3.11\,\rm mHz$, with an expected noise reduction by a factor of $1.8$. Instead, the central plot of Fig. 3 shows that the reduction was by a factor of $4$, due to the fact that in going to spin mode V3 the best fitting line given by Eq.(4) jumped downwards (together with all fluctuations around it). At the new signal frequency of $3.11\,\rm mHz$ the limiting noise is still a factor between 2 and 3 higher than expected (Touboul et al (2019), Fig. 11 top plot).
@@
It is apparent from Eq.(4) that the vertical position of the line (and of the plotted curve) is dictated by the acceleration $f_{TkQ}$, given by Eq.(5). The parameters that this acceleration depends upon can hardly have changed, in going from V2 to V3 mode, by the amount required to explain the extra gain factor of about 2.2, never mentioned in later publications and left unexplained to this date. Indeed, this event may not be exceptional, or due to the transition from lower to higher spin rate, since we observe very similar jumps even in sequential sessions at the same spin rate.
@@
Figure 3: SUEP session #218, Slide 24 of the talk given by M. Rodrigues at the 656th WE-Heraeus Seminar on “Fundamental Physics in Space” in Bremen, 2017. By increasing the spin frequency from mode V2 to mode V3 the frequency of the signal $\nu_{{}_{EP}}$ ($f_{{}_{EP}}$ in the slide) is moved to a higher value where noise is lower (as $1/\sqrt{\nu}$). This will ensure a noise reduction by a factor 1.8. However, it also happens that the best fitting line expressed by Eq.(4) undergoes a downward parallel translation (central plot) to a lower level of noise at all frequencies, including the new $\nu_{{}_{EP}}=3.11\,\rm mHz$, such that the total noise reduction is by a factor of 4 (unexplained).
@@
We cannot exclude a variation of the stiffness after a calibration, because in the calibration the quadratic acceleration term must be kept sufficiently small, which is achieved by offsetting the TM to a new zero, hence affecting the stiffness, $f_{{}_{TkQ}}$ and $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ (G. Catastini, private communication). Conversely, in sequential sessions, or within the same session analyzed with equally valid methods, we expect no such variations. Should any anomalies occur, they will clearly appear in the values of $Q$ because $Q\propto 1/S_{a}(\nu_{{}_{EP}})$. Moreover, the extent and the relevance of the anomaly can be assessed by comparison with $Q$ measurements carried out in ground tests of WEP as well as in experiments for the measurement of the universal constant of gravity and the detection of gravitational waves.
@@
In order to carry out this analysis we need the values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ for all the valid sessions, but there is no table listing these values in the Microscope publications, as there is for $\delta_{i}\pm\sigma_{i}$ (Touboul et al (2022b)) although accelerations are the measured physical quantities while $\delta$ is a derived number (see Touboul et al (2022b), Eq. (6)) and moreover $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ does not depend on the integration time –unlike $\delta$. With all $\delta$ values from Touboul et al (2022b), Tables 6, 7 and the duration of each session from Rodrigues et al (2022a) we use Eq.(6) to derive all the corresponding values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$. The correctness of our derived values can be checked when the measured ones are published. [^2] In Chhun et al (2022), Fig. 11, the values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ are plotted for all sessions, including those which have not contributed to the final result of the WEP test, the purpose being to demonstrate an overall stability after the first few sessions of the mission. Although the numerical values of the plotted marks are not easy to read in the vertical scale used, for the 120-orbit SUEP session #236 (V3 mode, blue curve in Fig. 11, error bar shown on V2 mark by mistake) we can see that the plotted mark is much too high (by at least a factor of 3) with respect to the best fit value of about $2\times 10^{-11}\,\rm ms^{-2}/\sqrt{Hz}$ reported in Fig. 6, Touboul et al (2022b). The discrepancy is confirmed using Eq.(6), because the plotted mark leads to a value of $\delta_{236}$ at least 3 times higher than reported in Touboul et al (2022b), Table 7. For other sessions of 120-orbit duration like #236, such as #238, #254 and #404 we know from Eq.(6) that the ratios of their $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ values to that of session #236 must be the same as the ratios of the corresponding values of $\delta$ (available in Table 7 of Touboul et al (2022b)). While these ratios are never close to 1, their individual marks are all plotted at about the same height in Fig. 11 of Chhun et al (2022). Thus, if this figure is correct, it contradicts Table 7 of Touboul et al (2022b) as well as Fig. 1 of Touboul et al (2022a) and the final Microscope test of WEP published therein.
@@
Then, given the equilibrium temperature (Touboul et al (2022b)), the masses and the measured stiffnesses, the quality factor $Q$ is inferred from Eq.(5). Since $k_{1}\,,k_{2}$ may vary because of calibrations we compute also $k/Q$ assuming the same unknown $k$ for both TM’s. These quantities are listed in Table 1 for SUEP and in Table 2 for SUREF, along with the percentage $p$ of artificial data that have been introduced in each session after the elimination of glitches.
@@
Table 1: SUEP Sensor Unit. Session number in V2 or V3 spin mode; session $\delta$ and percentage of glitches $p$ from Tables 6,7 and 4,5 of  Touboul et al (2022b); calculated values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$, $Q$, $k/Q$. $\downarrow$ indicates a jump between sequential sessions; * an $M$ vs $A$ discrepancy (see text below).
@@
Two methods of analysis, $M$-$ECM$ ($M$) and $ADAM$ ($A$), have been employed to estimate the experiment parameters, including $\delta$ and $\sigma$. Both methods use artificially reconstructed data to account for the missing data resulting from the elimination of glitches. [^3] Method $M$ estimates in the time domain the missing data, maximizing the likelihood conditional on the observed data, and then computes the least-squares estimate of the regression parameters. An estimation of the SD is also produced in the process. Method $A$ performs the parameter estimation in the frequency domain. As such it requires an uninterrupted, regularly spaced time series, which is obtained by filling the gaps left by the removal of glitches with the artificial data estimated by $M$. The two methods are stated to be equivalent, but in a few cases the values of $\delta$ calculated by $M$ and $A$ differ considerably (by a factor $\simeq 2$ in sessions #218, #438 and by a factor 7 in #442), implying that the same data stream leads to a different SD according to one or the other method. SUEP session #442 is an instructive extreme case. As reported in Touboul et al (2022b), Table 7, the noise is essentially the same but the values of $\delta$ differ by a factor 7; $\delta$ representing a level of WEP violation compatible with the measured acceleration difference noise. That is, for the same measurement session, in the same experimental conditions, the same data with the same level of noise lead to almost one order of magnitude difference in the evaluation of $\delta$. The difference is particularly evident when the results are interpreted in terms of the physics of the dominant thermal noise from internal damping (widely different values of $Q$, up to an unrealistic large value in session #442). SUREF session #778-1 is the only one, among all 32 sessions ($19$ SUEP $+$ $13$ SUREF), not affected by glitches, hence without artificial data. In this case, the results are exactly the same for $\delta$ and only slightly different for $\sigma$, as one expects if the same data set is analyzed with two equally valid methods (Touboul et al (2022b), Table 6). Sessions #442 and #778-1 show that the $M$ and $A$ methods are in agreement when applied to a time series of real measured data, but give different results (in particular different $\delta$ values) in the presence of artificial reconstructed data. Since the artificial data in $A$ are those estimated by $M$, it is the way the two methods manipulate these data that makes the difference.
Table 2: Same as Table 1 for SUREF Sensor Unit. A $\bullet$ has been added to indicate sessions with $Q$ values much larger than the largest value of $118$ measured in ground tests (Willemenot and Touboul (2000)).
@@
In SUEP session #218 the values of $\delta$ differ by a factor of about 2 depending on the type of analysis. This session (120 orbits) was the basis of the early results (Touboul et al (2017)) and was further elaborated two years later (Touboul et al (2019)), ending up with a different value of the spectral density at the signal frequency. Even discarding the earlier results, it is apparent that the latest analysis cannot be considered conclusive. $Q$ values differing by a factor 3.5, for the same oscillator in the same conditions, are inexplicable.
The reference sensor SUREF behaves quite differently from SUEP. Thermal noise is about a factor $\sqrt{2}$ smaller than for SUEP (assuming the same $Q$), due to a factor $2$ smaller $(k_{1}/m_{1}^{2}+k_{2}/m_{2}^{2})$ term, and residual systematic errors are comparable to the stochastic errors (Touboul et al (2022b)). Four out of nine sessions have been split because of sudden jumps in the mean value of the differential acceleration that required the two data segments to be treated as distinct experiments (in SUEP only one session, #326, out of 18 has been split). Unlike glitches, these jumps do not occur on all accelerometers simultaneously, hence they cannot be attributed to the spacecraft but are likely to originate in the accelerometers themselves, and last much longer (tens of seconds). In SUREF, in addition to the three cases of anomalous jumps mentioned above, we observe three sessions with large $M$ vs $A$ discrepancies (see Table 2). The values of $Q$ are typically larger than in SUEP and in three cases much larger than the largest value of $118$ reported in ground measurements (Willemenot and Touboul (2000)).
Early results (Touboul et al (2017)) were based for SUREF on session #176 (62 orbits). The spectral density $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ reported in Touboul et al (2017) was confirmed in Touboul et al (2019) and attributed to $1/\sqrt{\nu}$ thermal noise, yielding $\delta=3.75\times 10^{-15}$ and $Q=65$, consistent with ground measurements. Instead, the final values of $\delta_{{M}}$ and $\delta_{{A}}$ are much smaller (Table 2), and the corresponding $Q$’s of 317 and 283 are far too high to be realistic on the basis of ground tests. Note that session #176 has $40\%$ of reconstructed data.
@@
As shown in Sec. 3, the control of the TMs requires to control large acceleration biases. The acceleration bias $a_{bias}$ and the voltage $V_{bias}$ applied to control it are related through the scale factor $s_{{}_{f}}$:
@@
$s_{{}_{f}}=\frac{2\pi\epsilon_{\circ}(2r)}{md}\cdot(V_{p}-V_{p}^{\prime})$ (8)
@@
$\delta a{{}_{fpp}}=\frac{2\pi\epsilon_{\circ}(2r)}{md}\cdot\delta V_{fpp}\cdot V_{bias}\ \ \ \ .$ (9)
Using Eqs.(7) and (8), it leads to the relation
@@
As recalled in Sec. 3, the difference of acceleration biases between the TMs of SUEP is $\simeq 10^{-6}\,\rm ms^{-2}$. According to Eq.(10), a fluctuating patch potential $\delta V_{fpp}\simeq 0.3\,\mu V$ at the V3 signal frequency $\nu_{{}_{EP}}=3.11\,\rm mHz$ would produce a spurious acceleration of $4\times 10^{-14}\,\rm ms^{-2}$, as large as the average acceleration difference measured in all SUEP sessions (once normalized by session duration) at $\nu_{{}_{EP}}$ (Chhun et al (2022), Fig. 10) and enough to change it by $100\%$ either way.
In GOCE a source of low frequency acceleration noise with $1/\sqrt{\nu}$ dependence below $0.01\,\rm Hz$ was found to be due to fluctuating potentials generated by charge patches on the test mass and its electrodes (Catastini and Dumontel (2002)). GOCE was set up as a gravity gradiometer with $0.5\,\rm m$ separation of the TMs and a measurement bandwidth from $500\,\rm mHz$ to $0.1\,\rm Hz$. At $3\,\rm mHz$ the acceleration difference sensitivity was of $2\times 10^{-12}\,\rm ms^{-2}/\sqrt{Hz}$ along the ultra sensitive $Y,Z$ axes (more than one order of magnitude better than that of SUEP sensitive axis) and $4\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ along the least sensitive $X$ axis, with a contribution from patch fluctuating potentials of about $4\times 10^{-13}\,\rm ms^{-2}/\sqrt{Hz}$ and $2\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ respectively (Touboul et al (2016)). [^4] With no wire the zero of each TM control (dictated by geometry, construction, mounting and electronics) does not coincide with the zero of other forces acting on the test mass, such as Earth tides and electrostatic stiffness. Applying a constant bias is therefore inevitable in the Microscope design, regardless of the gold wire. Hence, any fluctuation of the bias in the frequency region of the WEP violation signal is relevant when aiming at very high precision.
@@
An unstable zero is a known issue in WEP tests, related to the problem of initial conditions (or release) errors, affecting all tests not designed as intrinsically null experiments, including those with laser tracked satellites, celestial bodies and cold atoms (Blaser (2001); Nobili et al (2008); Nobili (2016)). In other tests, primarily those based on torsion balances, the physical observable is a true null (Nobili and Anselmi (2018a)).
5 Outliers, gaps and artificial data